mathslover (mathslover):

Not getting this... \(I = \int \frac{1}{x^4 - 1} dx\)

mathslover (mathslover):

@vishweshshrimali5 i know that x^4-1 = (x^2+1) (x+1)(x-1)

mathslover (mathslover):

but, when we get to this : \[\large{I = \int \cfrac{1}{(x+1)(x-1)(x^2+1)} }\]

OpenStudy (isaiah.feynman):

I'd try partial fractions.

mathslover (mathslover):

\[ I =\cfrac{1}{2} \int \left[ \cfrac{1}{x-1} - \cfrac{1}{x+1} \right] \cfrac{1}{x^2+1} dx \]

OpenStudy (vishweshshrimali5):

No convert all of them into fractions like this: \[\large{\cfrac{1}{(x+1)(x-1)(x^2+1)} = \cfrac{A}{x+1} + \cfrac{B}{x-1} + \cfrac{Cx + D}{x^2+1}}\]

mathslover (mathslover):

This is really confusing : \((A+B+C)x^3 + (B-A+D)x^2 + (A+B-C) x + (B-A-D) = 1\)

mathslover (mathslover):

this has become more like an cubic equation. I want all the other terms except (B-A-D) to be zero, right? Can x be zero?

OpenStudy (vishweshshrimali5):

We are not putting values of x here. It is some general equation. So, try equating the powers of x both sides.

OpenStudy (vishweshshrimali5):

The coefficient of x^3 in LHS was A+B+C but in RHS it is 0. So, A+B+C = 0.

OpenStudy (vishweshshrimali5):

And so on...

mathslover (mathslover):

A + B + C = 0 B - A + D = 0 A + B - C = 0

mathslover (mathslover):

And B- A - D = 1

OpenStudy (vishweshshrimali5):

Yes

mathslover (mathslover):

Now, that is confusing. C = 0 A = -B 2B + D = 0 B = -D/2 B = A + D + 1 2B = D + 1 -D/2 = D + 1 3D/2 = -1 D = -2/3

mathslover (mathslover):

Since, A + B + C = 0 and A + B - C = 0 Clearly, A + B = C from 2nd So, 2C = 0 , and C = 0.

mathslover (mathslover):

A + B = 0 => A = -B ---------(1) B - A + D = 0 2B + D = 0 B = - D/2

mathslover (mathslover):

B = A + D + 1 Oh.. okay, got the mistake there. 2B = D + 1 -D = D + 1 D = -1/2

mathslover (mathslover):

C = 0 ; B = 1/4 ; A = -1/4 ; D = -1/2 ...

OpenStudy (vishweshshrimali5):

Perfect

OpenStudy (vishweshshrimali5):

Yes now you can do simple integration.

OpenStudy (mathmale):

I'd go along with this:\[\large{\cfrac{1}{(x+1)(x-1)(x^2+1)} = \cfrac{A}{x+1} + \cfrac{B}{x-1} + \cfrac{Cx + D}{x^2+1}}\] ... Looks like you're already caught up with the task of determining A, B, C and D.

mathslover (mathslover):

Okay, so that becomes : \(\cfrac{-1}{4(x+1)} + \cfrac{1}{4(x-1)} - \cfrac{1}{2(x^2 + 1)} \)

OpenStudy (mathmale):

Yes, assuming that your A, B, C and D are correct. Integrate term by term. In your shoes I'd write the first term as \[\frac{ -1 }{ 4 }\int\limits_{a}^{b}\frac{ dx }{ x+1 }.\]

OpenStudy (mathmale):

Or leave this integral indefinite and skip the limits of integration.

mathslover (mathslover):

So, now I have to use this : \[ \int \cfrac{1}{ax+ b } = \cfrac{1}{a} \log |ax+b| \] Right

OpenStudy (vishweshshrimali5):

Right

OpenStudy (mathmale):

You CAN use that formula, but you don't HAVE TO. Seeing the following:\[\frac{ -1 }{ 4 }\int\limits\limits_{a}^{b}\frac{ dx }{ x+1 }\]

OpenStudy (mathmale):

I would immediately write \[\frac{ -1 }{ 4 }\ln |x+1|+C\] ... but either approach should give you the correct integral.

mathslover (mathslover):

Oh, yes! Got it sir. That approach will be easier, I guess.

OpenStudy (vishweshshrimali5):

Yeah I would suggest you to do the same as @mathmale sir has suggested to do.

OpenStudy (mathmale):

\[\frac{ -1 }{ 4 }\ln |x+1|\]can be re-written as\[\ln |x+1|^{\frac{ -1 }{ 4 }}\] if desired, and the result either left as is or the expression manipulated so that you'd have a positive (instead of negative) exponent. There are other ways of doing this, as well.

mathslover (mathslover):

Okay, so, I will use this then : \(\ln (a) - \ln(b) = \ln \left( \cfrac{a}{b} \right) \)

OpenStudy (mathmale):

Wouldn't you want to use the law of logarithms that pertains to powers? The rule you've just typed in here pertains to quotients.

OpenStudy (mathmale):

It just might help to ignore the coefficients for now and determine, using only rules of logs, what the end result will look like. THEN, consider the coefficients and determine how they modify your previous result.